OpenStudy (anonymous):

area of a rectangle in simplest form 3 yards by 2 1/2 yards

OpenStudy (anonymous):

A = L W = (3)(5/2)

OpenStudy (anonymous):

3 ( 2+(1/2) ) = 15/2 = 7+(1/2)

OpenStudy (anonymous):

3x2 and 1/2 3 x2 = 6 3x 1/2 = 1.5 or 1 and 1/2 so 6 + 1.5 is 7.5